### v1

# Yogesh Kulkarni, Pooyan Fazli

Abstract
Multimodal reasoning over long-horizon video is challenging due to the need for precise spatiotemporal fusion and alignment across modalities. While recent methods such as Group Relative Policy Optimization (GRPO) have shown promise in this domain, they suffer from three key limitations: (1) data inefficiency from their on-policy design, (2) a vanishing advantage problem, where identical or near-identical rewards within a group eliminate the learning signal by producing zero-valued advantages, and (3) uniform credit assignment that fails to emphasize critical reasoning steps. We introduce AVATAR (Audio-Video Agent for Alignment and Reasoning), a framework that addresses these limitations through two core components: (1) an off-policy training architecture that improves sample efficiency and resolves vanishing advantages by reusing past experiences with greater reward diversity, and (2) Temporal Advantage Shaping (TAS), a novel credit assignment strategy that upweights key reasoning phases during learning. AVATAR achieves strong performance across various benchmarks, outperforming the Qwen2.5-Omni baseline by +5.4 on MMVU, +4.9 on OmniBench, and +4.5 on Video-Holmes, while demonstrating over 35% higher sample efficiency.
Introduction
Multimodal large language models (MLLMs) must align video, audio, and language modalities to support long-horizon reasoning (Lu et al. 2025). This requires balancing dense temporal coverage for narrative understanding with high spatial resolution for visual grounding (Zhong et al. 2025; Chen et al. 2025b). Reinforcement learning (RL), particularly Group Relative Policy Optimization (GRPO)(DeepSeek-AI et al. 2025; Chen et al. 2025a), has emerged as a promising approach for enhancing such reasoning. However, applying GRPO directly to multimodal domains reveals key limitations.
Figure 1: Standard GRPO (top) vs. AVATAR (bottom). AVATAR enhances standard GRPO with two key components: (1) an off-policy architecture using a stratified replay buffer to improve data efficiency, and (2) Temporal Advantage Shaping (TAS), a novel credit assignment strategy that focuses learning on critical reasoning steps.
First, GRPO is an on-policy method, discarding samples after each update. This makes it data-inefficient, a critical issue in the video domain where learning models rely on expensive, large-scale annotated data (Liu et al. 2025c; Wang et al. 2025b, e). Prior methods like RePO (Li et al. 2025c) reuse past data via replay buffers but lack selective control over replay quality. Second, GRPO suffers from the vanishing advantage problem: when the reward variance within a group collapses, such as in cases where all responses are either uniformly correct or uniformly incorrect, the resulting zero-valued advantage signals prevent effective learning (Wang et al. 2025a; Liu et al. 2025a; Dai, Yang, and Si 2025). Third, GRPO suffers from uniform credit assignment (Samineni et al. 2025; Deng et al. 2025), assigning the same reward to all tokens in a reasoning chain, regardless of their contribution to the final response (Dang et al. 2025; Fan et al. 2025a; Meng et al. 2025). This issue is amplified in multimodal tasks, where the GRPO objective fails to distinguish visually grounded reasoning from plausible text. By treating fluent language and accurate perception as equivalent, it enables the model to produce correct-sounding responses that ignore visual input, resulting in frequent perception errors (Wang et al. 2025d; Fan et al. 2025b).
To address these challenges, we introduce AVATAR (Audio-Video Agent for Alignment and Reasoning), a unified framework built on GRPO to enhance multimodal reasoning, with a focus on video tasks. To address GRPO’s data inefficiency, AVATAR adopts an off-policy training architecture that reuses past experiences through a stratified replay buffer. Samples are grouped based on how often the policy has succeeded on them in the past. This creates a curriculum that helps the model focus on samples it previously failed on, which is especially important for learning from sparse and challenging audio-video data. Moreover, to improve training stability and reward signal quality, we extend prior clue-grounded rewards (Lu et al. 2025; Lai et al. 2025; Huang et al. 2025) that check whether the model uses relevant audio or visual cues from the query. We introduce two new components: stepwise rewards, which give partial credit for intermediate reasoning, and a Video-Context Reference Score (VCRS), which uses a moving average of recent rewards for each query to maintain a stable and non-zero advantage signal. We also address the uniform credit assignment problem with a new method called Temporal Advantage Shaping (TAS). Unlike prior work that treats all tokens equally (Samineni et al. 2025) or depends on statistical variance (Dang et al. 2025), TAS guides learning based on a token’s position in the reasoning chain, placing greater weight on the beginning (planning) and end (synthesis) stages. This is especially useful for audio-video tasks, where accurate early grounding (e.g., locating the speaker) and final synthesis (e.g., combining speech and visual cues to identify the speaker) are critical. In the experiments, we apply AVATAR within a three-stage RL training pipeline to evaluate its effectiveness: starting with general visual reasoning, followed by audio-visual alignment, and concluding with fine-grained audio-based object localization. In summary, our primary contributions are:
  • We present AVATAR, an off-policy RL framework that enhances data efficiency in audio-video reasoning by leveraging past experiences through a difficulty-aware replay buffer, which focuses on challenging samples.
  • We introduce new approaches for RL reward signaling and credit assignment by giving partial credit to intermediate reasoning steps, stabilizing learning with a moving average of past rewards, and emphasizing the most critical parts of the reasoning process.
  • Experiments show that AVATAR achieves significant gains over the Qwen2.5 Omni baseline, with improvements of +5.4 on MMVU, +4.9 on OmniBench, and +4.5 on Video-Holmes, while demonstrating over 35% higher sample efficiency.
Related Work
Multimodal Understanding.
Recent MLLMs have made remarkable progress towards audio-video understanding (Bai et al. 2025; Zhu et al. 2025; Sun et al. 2025; Fu et al. 2024; Tang et al. 2025; Islam et al. 2025). Key efforts focus on architectural enhancements like dedicated audio branches and video connectors (Cheng et al. 2024), using clue aggregators to ground responses in question-specific details (Ye et al. 2024), or with token interleaving to enhance temporal understanding (Tang et al. 2025). However, these methods excel at direct video question answering but cannot rationalize their thoughts through reasoning traces. This limitation is problematic for complex audio-visual tasks requiring multi-step reasoning and fine-grained cross-modal alignment, such as localizing sound sources. In contrast, AVATAR enhances structured reasoning capabilities in audio-visual models using a sample-efficient RL approach.
Multimodal Understanding with RL.
Early multimodal efforts used preference optimization like DPO (Rafailov et al. 2023; Kulkarni and Fazli 2025a, b; Sarkar and Etemad 2025), but RL methods such as GRPO (DeepSeek-AI et al. 2025; Chen et al. 2025a) have shown greater promise for reasoning. Recent variants adapt GRPO to specific tasks: Video-R1 uses temporal contrastive rewards (Feng et al. 2025), HumanOmniV2 employs LLM-judged rewards for global audio-visual context (Zhao et al. 2025a), Omni-R1 targets audio-visual segmentation (Zhong et al. 2025), and GRIT generates spatial reasoning chains (Fan et al. 2025b). However, these approaches suffer from uniform credit assignment and sample inefficiency. AVATAR addresses these limitations through its off-policy architecture and TAS credit assignment. For audio-visual tasks, it’s U-shaped temporal weighting emphasizes early grounding (localizing sounds) and late synthesis (combining auditory and visual evidence), while the stratified replay buffer handles sparse cross-modal data effectively.
Preliminaries
Group Relative Policy Optimization
GRPO is an RL framework that can be used to fine-tune MLLMs by comparing multiple candidate responses generated by a behavior policy $\pi_{\theta_{\text{old}}}$ for a given prompt $q$ (DeepSeek-AI et al. 2025). For each group of $K$ candidate responses $\{o_{i}\}_{i=1}^{K}$, scalar rewards $R(o_{i})$ from predefined reward functions are assigned to each response. These rewards are then normalized to compute relative advantages $A_{i}$ within the group.
$A_{i}=\frac{R(o_{i})-\mu_{R}}{\sigma_{R}+\epsilon_{\text{adv}}},$ (1)
where $\mu_{R}$ and $\sigma_{R}$ are the mean and standard deviation of the group rewards $\{R(o_{i})\}_{i=1}^{K}$ and $\epsilon_{adv}$ is a small constant for numerical stability. These group-relative advantages guide the update of the target policy $\pi_{\theta}$, encouraging it to assign higher probability to relatively better responses. GRPO maximizes the following clipped surrogate objective with KL regularization (Li et al. 2025c):
$\begin{split}\mathcal{J}_{\text{GRPO}}(\theta)&=\mathbb{E}_{q\sim\mathcal{D},\{o_{i}\}\sim\pi_{\theta_{old}}}\left[\frac{1}{K}\sum_{i=1}^{K}\min(r_{i}A_{i},\right.\\ &\quad\left.\text{clip}(r_{i},1-\epsilon,1+\epsilon)A_{i})-\beta D_{KL}(\pi_{\theta}\|\pi_{\text{ref}})\right],\end{split}$ (2)
where $q$ is a prompt sampled from the dataset $\mathcal{D}$, $r_{i}$ is the importance sampling ratio $\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\text{old}}}(o_{i}|q)}$, $\epsilon$ controls the clipping range, and $\beta$ weights the KL divergence term to ensure stability by regularizing the updated policy $\pi_{\theta}$ against a reference policy $\pi_{\text{ref}}$.
Data Inefficiency Problem.
GRPO is an on-policy method that updates using only fresh rollouts from the current policy and discards past data. This limits sample reuse and leads to inefficiency, especially in video understanding tasks where data is costly to collect. Off-policy updates offer a more efficient alternative by reusing past experiences to reduce training costs.
Vanishing Advantage Problem.
A fundamental limitation of GRPO arises from its advantage estimation mechanism. When the rewards $R(o_{i})$ for all outputs in a group are identical or nearly identical (e.g., all zero for a particularly difficult prompt or all one for an easy prompt), the numerator of the advantage function, $R(o_{i})-\mu_{R}$, becomes zero for all samples $i$. This results in zero advantages $(A_{i}=0)$, eliminating the learning signal and preventing the policy from updating. Consequently, GRPO fails to learn from such training samples, a phenomenon referred to as the vanishing advantage problem (Wang et al. 2025a; Li et al. 2025c).
Credit Assignment Problem.
Another key limitation of GRPO is its overly simplistic credit assignment. In Equation 2, a single scalar advantage $A_{i}$ is computed for the entire output $o_{i}$ and applied uniformly to all tokens during policy updates. This ignores the varying importance of different reasoning steps. Tokens critical to initial planning receive the same learning signal as those from less important phases, failing to address the temporal credit assignment challenge that RL is designed to solve.
 [Uncaptioned image] AVATAR
We introduce AVATAR, an audio-video agent for multimodal alignment and reasoning, designed to address key limitations of GRPO. To tackle the vanishing advantage problem and resulting data inefficiency, AVATAR employs an off-policy training architecture using a stratified replay buffer. To overcome uniform credit assignment, it introduces Temporal Advantage Shaping (TAS), which modulates advantages (i.e., re-weighs the learning signal for each token) across the reasoning sequence. The following sections detail these components.
Algorithm 1 AVATAR
1:Input: Initial policy $\pi_{\theta_{init}}$, Reward functions $R$, Hyperparameters $\alpha,\lambda_{\text{TAS}},\beta,\epsilon,K_{on},K_{off}$.
2:Initialize: Policy $\pi_{\theta}\leftarrow\pi_{\theta_{init}}$, Reference policy $\pi_{\text{ref}}\leftarrow\pi_{\theta_{init}}$, empty Difficulty-Aware Buffer $\mathcal{B}$.
3:for each training step do
4:  Sample a batch of prompts $Q_{batch}$ from dataset $\mathcal{D}$.
5:  $\pi_{\theta_{old}}\leftarrow\pi_{\theta}$ {Set behavior policy}
6:  {Experience Generation}
7:  Generate $K_{on}$ on-policy experiences $E_{on}=\{(q,o,R(o))\}$ using $\pi_{\theta_{old}}$ and compute advantages $A_{on}$.
8:  Sample $K_{off}$ off-policy experiences $E_{off}=\{(q,o,\pi_{\theta_{off}})\}$ from $\mathcal{B}$ and compute advantages $A_{off}$.
9:  {Temporal Advantage Shaping (TAS)}
10:  for each sample $o_{i}$ with advantage $A_{i}$ in $E_{on}\cup E_{off}$ do
11:   Compute per-token weights $w_{i,t}$ using Equation 6.
12:   Create token-level shaped advantage: $A^{\text{TAS}}_{i,t}\leftarrow A_{i}\cdot w_{i,t}$.
13:  end for
14:  {Policy Update}
15:  Compute total loss $\mathcal{J}_{\text{total}}(\theta)$ by combining the on-policy and off-policy surrogate objectives using their respective shaped advantages ($A^{\text{TAS}}_{on},A^{\text{TAS}}_{off}$) and the final KL penalty term.
16:  Update policy: $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{J}_{\text{total}}(\theta)$.
17:  {Buffer Update}
18:  Update Buffer $\mathcal{B}$ with the new on-policy experiences.
19:end for
Off-Policy Architecture
To address both the data inefficiency of on-policy learning and the vanishing advantage problem, AVATAR employs an off-policy architecture.
Replay Buffer.
Our framework uses a stratified replay buffer $\mathcal{B}$ divided into three dynamic tiers: Easy (25%), Medium (35%), and Hard (40%). Experiences are assigned based on the policy’s moving average reward $\bar{R}(q)$ for each prompt $q$, forming a progressive learning curriculum. Fixed tier capacities ensure rapid turnover of easy samples while retaining harder ones for repeated training.
Buffer Update and Hinting.
The buffer uses two metrics: the moving average reward $\bar{R}(q)$ to assign difficulty tiers, and KL-divergence $D_{KL}(\pi_{\theta}||\pi_{\beta})$ between the target policy and behavior policy to monitor policy stability. If a prompt remains hard (low $\bar{R}(q)$) and the policy stops exploring (low KL), a hint mechanism is triggered. Pre-computed hints are generated by providing the full problem context (video, audio caption, query, and ground truth) to Qwen2.5-VL (Bai et al. 2025) (e.g., “first locate the object making the sound, then count”) and are added to guide learning and escape local optima. This keeps training in a “challenging but solvable” zone, applying curriculum learning principles while enabling targeted exploration (Huang et al. 2025).
Off-Policy Update with Selective Sample Replay.
At each training step, we form a batch using both newly generated on-policy samples and off-policy samples from the replay buffer. To handle policy drift in off-policy data, we apply importance sampling. The full objective function for AVATAR combines the standard on-policy GRPO loss with a corrected off-policy term:
$\mathcal{J}_{\text{AVATAR}}=\mathcal{J}_{\text{on-policy}}(\theta)+\alpha\cdot\mathcal{J}_{\text{off-policy}}(\theta),$ (3)
Where $\alpha$ is a hyperparameter balancing the off-policy contribution, and the off-policy objective $\mathcal{J}{\text{off-policy}}$ follows Equation 2, with the importance sampling ratio $r_{i}$ computed using the generating policy $\pi{\theta_{\text{off}}}$ stored in the replay buffer:
$r_{i}^{\text{off}}(\theta)=\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{off}}(o_{i}|q)}.$ (4)
This hybrid training strategy, inspired by recent work on off-policy PPO and GRPO (Li et al. 2025c; Mroueh et al. 2025), enables AVATAR to be significantly more data-efficient, while the buffer ensures that the replayed samples provide a stable and informative learning signal. The training loop is detailed in Algorithm 1.
Temporal Advantage Shaping (TAS)
Solving complex tasks typically requires more reasoning effort during the initial planning and final synthesis stages than during intermediate evidence gathering. To address GRPO’s uniform credit assignment issue, we introduce Temporal Advantage Shaping (TAS), a method that adjusts each token’s learning signal based on its position in the reasoning chain. This refines the single-advantage formulation used in standard GRPO (Equation 2). As shown in Figure 2, TAS applies a U-shaped weighting curve that amplifies the contribution of tokens in the planning and synthesis phases, unlike the flat, uniform credit assignment used in standard GRPO. This prevents overemphasizing low-value tokens and better reinforces critical reasoning steps.
Figure 2: To address GRPO’s uniform credit assignment (gray line), TAS applies a U-shaped weight to amplify advantages during crucial planning and synthesis stages.
U-Shaped Weight Calculation.
For a reasoning sequence of length $L$, we normalize each token’s position $t\in\{0,1,\ldots,L-1\}$ to the range $[0,1]$ by defining
$\tilde{t}=\frac{t}{L-1}.$ (5)
We then apply a parabolic function to generate a U-shaped weight $w_{t}$:
$w_{t}=1.0+\lambda_{\text{TAS}}\cdot(2\tilde{t}-1)^{2},$ (6)
where $\lambda_{\text{TAS}}$ controls the curve’s amplitude. This function assigns the minimum weight of 1.0 to tokens at the midpoint of the sequence ($\tilde{t}=0.5$) and the maximum weight of $1.0+\lambda_{\text{TAS}}$ to tokens at the start and end ($\tilde{t}=0$ or $1$). These weights produce a token-specific shaped advantage: $A^{\text{TAS}}_{i,t}=w_{i,t}\cdot A_{i}$ where $i\in\{1,\dots,K\}$ indexes each of the $K$ candidate responses in the group. AVATAR’s loss is formed by replacing the standard GRPO advantage $A_{i}$ in Equation 2 with this shaped advantage.
A key challenge is that reasoning sequence lengths, $L_{i}$, vary across completions. Our TAS mechanism is robust to this variability because the normalization step in Equation 5 makes the weight for any token $t$ depend on its relative position $\tilde{t}$ within the sequence, rather than its absolute position. This length-agnostic approach is supported by our ablation studies (Table 4), which show that our U-shaped weighting consistently outperforms the uniform baseline.
Figure 3: Three-stage RL training pipeline to evaluate AVATAR. The framework advances from Cold start SFT (Stage 0) to Visual Reasoning (Stage 1) to Audio-Visual Reasoning (Stage 2) to Audio-Object Localization (Stage 3).
Reward Functions
AVATAR is guided by a set of reward functions as detailed below:
(1) Format Reward ($R_{\text{format}}$).
A binary reward that verifies adherence to the predefined reasoning format (i.e., <think>...</think><answer>...</answer>).
$R_{\text{format}}(o_{i})=\begin{cases}1,&\text{if format is correct}\\ -1,&\text{otherwise}\end{cases}$ (7)
(2) Final Answer Accuracy ($R_{\text{acc}}$).
A black-box evaluation of the final answer $a_{i}$ extracted from the <answer> tag, compared to the ground truth $a_{\text{GT}}$. For numerical tasks, we use the relative Mean Absolute Error (rMAE) for a dense reward signal (Lu et al. 2025).
$R_{\text{acc}}(a_{i})=1.0-\min\left(1.0,\frac{|a_{i}-a_{\text{GT}}|}{a_{\text{GT}}}\right)$ (8)
(3) Self-Rewarding ($R_{\text{self}}$).
We use self-rewarding to help the model learn from its own consensus. Given a group of $K$ generated answers $\{o_{1},o_{2},\dots,o_{K}\}$ for a prompt, we choose the most common answer as a “pseudo-correct” $o^{*}_{\text{maj}}$ one using majority vote.
$o^{*}_{\text{maj}}=\text{argmax}_{o\in\{o_{1},\dots,o_{K}\}}\sum_{i=1}^{K}\mathbb{I}(o_{i}=o),$ (9)
where $\mathbb{I}$ is the indicator function. The self-reward for each answer $o_{i}$ is then a binary score based on its agreement with this consensus:
$R_{\text{self}}(o_{i})=\begin{cases}1,&\text{if }o_{i}=o^{*}_{\text{maj}}\\ 0,&\text{otherwise}\end{cases}$ (10)
(4) Stepwise Reasoning Judge ($R_{\text{judge}}$).
A white-box reward from a frozen VLM judge (InternVL-3B (Zhu et al. 2025)), which provides a detailed score for reasoning quality within the <think> block. It evaluates logical consistency and correct use of grounding clues from the prompt. The score $s_{\text{judge}}\in[0,1]$ reflects the reasoning process quality independently of the final answer.
Video-Context Reference Score (VCRS).
To ensure a stable learning signal, we introduce the Video-Context Reference Score (VCRS), named for its role as a stable reference score based on the historical context of a specific video prompt. VCRS acts as a multiplicative factor in the final advantage calculation. It is computed as a moving average of rewards over the last 20 processed instances of a given prompt, ensuring the advantage baseline is never zero and helping to prevent learning stagnation on difficult tasks.
We apply AVATAR within a three-stage RL training pipeline (Figure 3) to evaluate its effectiveness. Stage 1 develops general visual reasoning, Stage 2 introduces audio-visual alignment, and Stage 3 addresses fine-grained audio-based object localization. Each stage leverages specific reward functions to ensure a stable and informative learning signal: Stage 1 uses format and accuracy rewards, Stage 2 augments these with the self-rewarding mechanism, and Stage 3 combines format rewards with the stepwise reasoning judge.
Audio-Visual Benchmarks
Model Reasoning OmniBench DailyOmni AV-Counting AV-Odyssey WorldSense IntentBench
Qwen2.5 Omni $\times$ 44.2 44.0 22.3 29.8 44.2 63.7
Echolnk $\checkmark$ 46.5 46.2 22.7 31.1 45.7 63.6
Omni-R1 $\checkmark$ 46.9 46.8 22.0 31.2 44.1 63.5
HumanOmni $\checkmark$ 44.9 47.6 19.6 30.3 45.4 *
Ola-7B $\times$ 45.3 52.3 17.4 25.6 44.2 59.1
AV-Reasoner $\checkmark$ 48.3 53.8 23.0 25.6 44.6 59.5
AVATAR $\checkmark$ 49.1+4.9 47.0+3.0 23.1+0.8 32.1+2.3 46.0+1.8 63.9+0.2
Table 1: Comprehensive evaluation of AVATAR vs. state-of-the-art audio-video understanding models. Best scores are in bold, and second-best scores are underlined. For AVATAR, performance improvements over Qwen-2.5-Omni are shown as subscripts. All results are reproduced by us and * refers to data contaminated.
General Video Understanding Video Reasoning
Model Reasoning MVBench Video-MME LVBench Video-Holmes MMVU TOMATO
Omnimodal Baselines (with Audio)
Qwen2.5 Omni $\times$ 66.1 58.3 37.2 40.6 60.2 29.0
Echolnk $\checkmark$ 66.2 60.8 37.6 42.5 65.7 29.9
Omni-R1 $\checkmark$ 66.0 60.7 37.6 44.2 63.6 29.2
HumanOmni $\checkmark$ 61.4 63.1 36.2 39.6 61.8 27.1
Ola - 7B $\times$ 40.1 59.1 35.5 40.1 56.6 -
AV-Reasoner $\checkmark$ 47.9 56.8 33.7 39.6 57.9 -
Video-only Reasoning Baselines (on Qwen2.5-VL)
Qwen2.5-VL $\times$ 59.4 60.1 31.2 38.0 61.9 29.3
Video-R1 $\checkmark$ 63.9 59.3 27.8 41.0 63.1 19.8
VideoRFT $\checkmark$ 62.1 59.8 34.2 40.5 58.0 20.6
TW-GRPO $\checkmark$ 63.3 60.1 33.8 38.4 61.4 23.5
AVATAR $\checkmark$ 66.4+0.3 62.8+4.5 38.4+1.2 45.1+4.5 65.6+5.4 30.8+1.8
Table 2: Comprehensive evaluation of AVATAR on general video understanding and reasoning benchmarks. Best scores are in bold, and second-best scores are underlined. For AVATAR, performance improvements over Qwen-2.5-Omni are shown as subscripts. All results are reproduced by us.
Model Configuration Audio-Visual Video Reasoning
OmniBench DailyOmni AV-Odyssey Video-Holmes MMVU TOMATO
(1) Baseline Qwen2.5 Omni 44.2 44.0 29.8 40.6 60.2 29.0
(2) Baseline GRPO 45.4+1.2 44.8+0.8 31.3+1.5 43.2+2.6 64.0+3.8 29.2+0.2
+ TAS Only 45.1+0.9 45.4+1.4 31.4+1.6 43.4+2.8 64.3+4.1 29.5+0.5
+ Curriculum & Replay 47.8+3.6 45.9+1.9 31.6+1.8 43.7+3.1 64.8+4.6 29.3+0.3
AVATAR (w/o Hinting) 48.2+4.0 46.1+2.1 31.8+2.0 44.3+3.7 65.0+4.8 30.2+1.2
AVATAR (w/ Hinting) 49.1+4.9 47.0+3.0 32.1+2.3 45.1+4.5 65.6+5.4 30.8+1.8
Table 3: Component-wise ablation study across audio-visual and video reasoning benchmarks.
Experiments
Implementation Details.
We use Qwen2.5-Omni-7B (Xu et al. 2025) as the backbone model and implement AVATAR using the MS-Swift (Zhao et al. 2025b) framework for GRPO training. As shown in Figure 3, we begin with a cold-start SFT phase on general image and video QA datasets, including Video-R1-CoT (Feng et al. 2025) and TPO (Li et al. 2025b). This is followed by three RL stages of increasing complexity: Stage 1 builds foundational visual reasoning using general video QA data from Video-R1 (Feng et al. 2025). Stage 2 incorporates audio-visual QA using AVQA (Yang et al. 2022) and AVE (Tian et al. 2018), with audio cues annotated via Kimi-Audio (KimiTeam et al. 2025). Stage 3 focuses on audio-visual object localization, using AVSBench (Zhou et al. 2022) and stepwise rewards from the InternVL-3B (Zhu et al. 2025) judge. To ensure computational efficiency while preserving temporal information, we uniformly sample 32 video frames during both training and evaluation. We also apply TAS throughout all RL stages. All experiments are conducted on 8 NVIDIA L40S GPUs.
State-of-the-art-models.
We compare AVATAR with a broad range of models across audio-visual and video-only modalities. For audio-visual models, we evaluate against Qwen2.5-Omni-7B (Xu et al. 2025), EchoInk (Xing et al. 2025), Omni-R1 (Zhong et al. 2025), HumanOmni (Zhao et al. 2025a), Ola-7B (Liu et al. 2025b), and AV-Reasoner (Lu et al. 2025). To assess the impact of audio integration, we also compare with SOTA video-only models: Qwen2.5VL (Bai et al. 2025), Video-R1 (Feng et al. 2025), Video-RFT (Wang et al. 2025b), and TW-GRPO (Dang et al. 2025).
Benchmarks.
For audio-visual tasks, we use OmniBench (Li et al. 2025e) (image-audio-text reasoning), DailyOmni (Zhou, Wang, and Wu 2025) (ambient sound scenarios), AV-Counting (Lu et al. 2025) (cross-modal counting), AV-Odyssey (Gong et al. 2024) (audio-visual alignment), WorldSense (Hong et al. 2025) (omnimodal understanding in-the-wild), and IntentBench (Zhao et al. 2025a) (reasoning about human intent and actions). For video understanding and reasoning, we evaluate on MVBench (Li et al. 2024) (temporal perception-to-cognition QA), Video-MME (Fu et al. 2025) (full-spectrum video comprehension), LVBench (Wang et al. 2024) (long-video reasoning), Video-Holmes (Cheng et al. 2025) (causal inference in suspense clips), MMVU (Zhao et al. 2025c) (domain-expert knowledge analysis), and TOMATO (Shangguan et al. 2025) (temporal reasoning).
Ablation Configuration Audio-Visual Video Reasoning
Group Setting OmniBench DailyOmni AV-Odyssey Video-Holmes MMVU TOMATO
Baseline Qwen2.5 Omni 44.2 44.0 29.8 40.6 60.2 29.0
Curriculum SFT Only 45.8+1.6 45.2+1.2 30.1+0.3 41.8+1.2 62.1+1.9 28.9-0.1
SFT + Stage 1 RL 47.2+3.0 45.8+1.8 30.6+0.8 43.5+2.9 64.2+4.0 29.4+0.4
SFT + Stages 1+2 RL 48.6+4.4 46.7+2.7 31.8+2.0 44.2+3.6 65.1+4.9 30.1+1.1
SFT + Stages 1+2+3 RL (Ours) 49.1+4.9 47.0+3.0 32.1+2.3 45.1+4.5 65.6+5.4 30.8+1.8
TAS Linear Decay Weights 48.3+4.1 46.2+2.2 31.6+1.8 44.3+3.7 64.8+4.6 30.2+1.2
Linear Incline Weights 48.5+4.3 46.4+2.4 31.7+1.9 44.6+4.0 65.0+4.8 30.4+1.4
No TAS (Uniform Weights) 47.8+3.6 45.9+1.9 31.6+1.8 43.7+3.1 64.8+4.6 29.3+0.3
U-Shaped Weights (Ours) 49.1+4.9 47.0+3.0 32.1+2.3 45.1+4.5 65.6+5.4 30.8+1.8
Table 4: Ablation studies on curriculum design, reward strategy, and TAS.
Results
Audio-Visual Reasoning.
Table 1 shows that AVATAR significantly outperforms the Qwen2.5-Omni baseline on audio-visual reasoning tasks, with absolute gains of +4.9 on OmniBench, +3.0 on DailyOmni, and +2.3 on AV-Odyssey. Unlike standard GRPO, which often stalls on overly easy or hard samples (due to vanishing advantages), our stratified replay buffer keeps training in a “challenging but solvable” zone. This mitigates overfitting and builds a more robust reasoning foundation, enabling AVATAR to outperform other GRPO-based methods, surpassing AV-Reasoner by 0.8 and Omni-R1 by 2.2 on OmniBench alignment tasks. TAS further enhances performance on benchmarks that require precise temporal alignment, such as AV-Odyssey. By creating larger, more targeted gradients for the initial tokens that identify an audio cue and the final tokens that synthesize the corresponding visual event, TAS directly reinforces the cross-modal causal links required for such tasks.
This targeted credit assignment enhances AVATAR’s performance on audio-visual tasks like AV-Counting, contributing to stronger generalization across diverse, real-world scenarios, as seen in WorldSense.
General Video Understanding and Reasoning.
As shown in Table 2, AVATAR achieves SOTA performance on general video understanding benchmarks, such as MVBench (66.4) and LVBench (38.4), and excels in complex reasoning tasks. In particular, it delivers significant gains on multi-step causal inference benchmarks, such as Video-Holmes (+4.5) and MMVU (+5.4), enabled by two key components: the stratified replay buffer, which ensures exposure to challenging data, and TAS, which guides reasoning by focusing on the planning and synthesis phases of the token sequence. This is especially evident when compared to other credit assignment methods. TW-GRPO’s statistical token weighting results in unstable performance and even degrades on temporal fine-grained tasks, such as TOMATO. In contrast, TAS leverages the reasoning chain’s temporal structure to provide a more stable learning signal, enabling AVATAR to consistently achieve top performance across reasoning benchmarks.
Ablation Studies
How efficient is AVATAR compared to baseline GRPO?
Figure 4 shows AVATAR’s superior sample efficiency and final performance over the GRPO baseline. GRPO plateaus at a 0.4 reward, solving only simpler queries due to uniform credit assignment. In contrast, AVATAR reaches 0.75 in just 2,500 iterations, a 35% efficiency gain. Its initial dip reflects the off-policy design, where the stratified buffer introduces harder samples. AVATAR quickly adapts via TAS-based credit assignment, converging efficiently while maintaining exploration at high reward levels. This shows its ability to learn from challenging samples that the baseline fails to leverage.
(a) Baseline Method
(b) AVATAR
Figure 4: Comparison of accuracy rewards during training.
What is the effect of each component in AVATAR?
Our ablation study (Table 3) highlights the impact of each component. Adding the replay buffer to standard GRPO yields substantial gains (e.g., +3.6 on OmniBench) by improving data efficiency and generalization. TAS alone also enhances performance (e.g., +4.1 on MMVU) by addressing uniform credit assignment and strengthening reasoning. Combining both further improves results, as the replay buffer enhances sample quality while TAS guides learning. Finally, the hint mechanism provides additional gains (e.g., +0.9 on OmniBench) by helping the model escape local optima, leading to the best overall performance of the full AVATAR system.
What is the effect of the curriculum and TAS shape?
Our ablations in Table 4 validate the effectiveness of both the three-stage training curriculum and the U-shaped TAS weighting. The curriculum shows a clear trend of consistently positive returns. Stage 1 RL yields the largest improvements over SFT (e.g., +2.1 on MMVU) by building core visual reasoning skills. Stage 2 adds further improvements (e.g., +1.4 on OmniBench) as the model learns to handle more complex cross-modal alignment, including audio cues. Stage 3 offers additional gains (e.g., +0.9 on Video-Holmes), reflecting the inherent difficulty of fine-grained localization tasks. The TAS ablation supports our core hypothesis: weighting only the planning (linear decay) or synthesis (linear incline) phases leads to weaker performance. In contrast, the U-shaped weighting, which emphasizes both early planning and late synthesis, proves most effective.
Config Halluc TP AC
Baseline GRPO 35.6 53.6 32.5
AVATAR w/o Off-Policy 40.1+4.5 54.8+1.2 34.1+1.6
AVATAR w/o TAS 48.9+13.3 56.1+2.5 35.8+3.3
AVATAR (Full) 51.2+15.6 57.2+3.6 37.3+4.8
Table 5: Ablation on key reasoning tasks from the WorldSense benchmark. Halluc: Hallucination, TP: Temporal Prediction, AC: Audio Change
How do AVATAR’s components impact audio-visual reasoning?
The impact is evident on reasoning tasks from the WorldSense benchmark (Table 5). The highest improvement is on the Hallucination task, a gain primarily driven by our off-policy architecture, which builds a more robust and grounded model by forcing repeated engagement with difficult samples from the hard tier. The effect of TAS is most clear on tasks requiring temporal logic, such as Temporal Prediction, which involves forecasting future events, and Audio Change, which requires identifying discrete shifts in the audio stream. The U-shaped weighting of TAS is uniquely suited for these problems, by modifying credit for initial tokens, it forces accurate grounding of the video’s initial state, while modifying credit for final tokens rewards correct synthesis of how that state evolved over time.
Conclusion
We introduce AVATAR, an RL framework that addresses key limitations of GRPO in multimodal audio-video reasoning. Our primary finding is that jointly addressing data inefficiency and uniform credit assignment is essential for robust performance. We show that an off-policy curriculum with a stratified replay buffer outperforms standard on-policy methods, while TAS significantly enhances reasoning quality by emphasizing critical planning and synthesis phases.
Acknowledgments
This research was supported by the National Eye Institute (NEI) of the National Institutes of Health (NIH) under award number R01EY034562. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
Appendix A Appendix
GRPO for Multimodal Reasoning
Recent works increasingly use RL methods, particularly GRPO (Xiao et al. 2025; Wan et al. 2025; Wang et al. 2025d), introducing new challenges related to training efficiency and stability. Wei et al. propose a “cold start” strategy via Supervised Fine-Tuning (SFT) before RL (Wei et al. 2025), while other approaches improve data efficiency through direct RL with test-time scaling  (Wang et al. 2025e). Concurrent advances enhance GRPO through curriculum learning (Lu et al. 2025), variance-aware data selection (Li et al. 2025a), hint-based guidance for difficult samples (Huang et al. 2025), and combining GRPO and DPO (Li et al. 2025d). Key limitations remain, including the “overthinking” issue, addressed by regulating reasoning length (Wang et al. 2025c; Xiao et al. 2025), and perception errors mitigated via grounded reasoning approaches (Fan et al. 2025b; Sarch et al. 2025) and new RL objectives such as PAPO’s Implicit Perception Loss (Wang et al. 2025d).
Audio-visual reasoning that incorporates audio often involves trade-offs between temporal coverage and spatial resolution. These challenges have been addressed through two-system architectures (Zhong et al. 2025), explicit contextual rewards (Zhao et al. 2025a), and task-specific improvements (Lu et al. 2025; Xing et al. 2025). AVATAR introduces a unified framework that integrates a stratified replay buffer for curriculum learning with Temporal Advantage Shaping (TAS) for credit assignment. This approach improves reasoning and grounding within a single model, eliminating the need for two-system architectures (Zhong et al. 2025) or explicit visual coordinate generation (Sarch et al. 2025).
(a) OmniBench
(b) DailyOmni
(c) AV-Odyssey
(d) Video-Holmes
(e) MMVU
(f) TOMATO
Figure 5: Performance vs. on-policy/off-policy sample ratio across six benchmarks. AVATAR achieves optimal performance with 4-4 split (4 on-policy, 4 off-policy samples), demonstrating that balanced mixing prevents both policy drift from excessive off-policy data and sample inefficiency from pure on-policy training.
Ablation Configuration Audio-Visual Benchmarks Video Reasoning Benchmarks
Group Setting OmniBench DailyOmni AV-Odyssey Video-Holmes MMVU TOMATO
Baseline Qwen2.5 Omni 44.2 44.0 29.8 40.6 60.2 29.0
Reward Suite $R_{\text{acc}}+R_{\text{format}}$ 46.8+2.6 45.2+1.2 30.1+0.3 42.3+1.7 62.1+1.9 28.9-0.1
$R_{\text{acc}}+R_{\text{format}}+R_{\text{self}}$ 47.5+3.3 45.8+1.8 30.7+0.9 43.1+2.5 63.4+3.2 29.4+0.4
$R_{\text{acc}}+R_{\text{format}}+R_{\text{self}}+R_{\text{judge}}$ 48.3+4.1 46.4+2.4 31.4+1.6 44.2+3.6 64.7+4.5 30.1+1.1
$R_{\text{acc}}+R_{\text{format}}+R_{\text{self}}+R_{\text{judge}}+VCRS$ 49.1+4.9 47.0+3.0 32.1+2.3 45.1+4.5 65.6+5.4 30.8+1.8
Table 6: Ablation studies on our reward strategy across three training stages. Stage 1 uses $R_{\text{acc}}$ and $R_{\text{format}}$ for basic accuracy and format compliance. Stage 2 adds $R_{\text{self}}$ (majority vote consensus) to the existing rewards, enabling consistency learning. Stage 3 incorporates $R_{\text{judge}}$ (stepwise VLM evaluation) alongside all previous rewards, providing detailed reasoning feedback. VCRS maintains stable advantage baselines through moving averages, preventing vanishing advantages. Each cumulative addition demonstrates clear performance gains, with largest improvements on reasoning-heavy benchmarks (Video-Holmes: +4.5, MMVU: +5.4) where structured feedback proves most beneficial.
Ablation Studies
What is the impact of the on-policy/off-policy sample ratio?
To determine the optimal balance between on-policy and off-policy data, we analyze performance across different sample ratios within a fixed group size of eight (Figure 5). We observe a consistent trend across all benchmarks: performance peaks when using a balanced 4-4 split (4 on-policy, 4 off-policy samples). Using fewer off-policy samples (0-3) fails to fully mitigate the data inefficiency of a purely on-policy approach. On the other hand, over-reliance on off-policy samples (5-7) degrades performance due to instability, as the policy drifts too far from the older, “stale” data in the replay buffer. The 4-4 split therefore represents an empirical sweet spot for AVATAR, maximizing the data efficiency gains of our off-policy architecture while maintaining the training stability necessary for effective learning.
(a) OmniBench
(b) DailyOmni
(c) AV-Counting
(d) AV-Odyssey
(e) WorldSense
(f) IntentBench
(g) MVBench
(h) Video-MME
(i) LVBench
(j) Video-Holmes
(k) MMVU
(l) TOMATO
Figure 6: Training curves across audio-visual and video reasoning benchmarks. AVATAR demonstrates superior sample efficiency and final performance, particularly on challenging reasoning tasks. AVATAR’s initial dip, followed by a strong recovery, validates the effectiveness of our off-policy architecture and TAS for credit assignment.
How do isolating rewards impact performance?
Table 6 shows the cumulative impact of each reward component in AVATAR’s training pipeline. The baseline, which includes only accuracy and format rewards, offers limited learning signals, especially on complex reasoning tasks such as Video-Holmes and MMVU, where binary feedback proves insufficient. Introducing self-rewarding ($R_{\text{self}}$) in Stage 2 enables consensus-based learning from the model’s own outputs, yielding moderate improvements on audio-visual tasks that benefit from internal consistency signals (e.g., +0.7 on OmniBench, +1.3 on MMVU). Adding the stepwise judge reward ($R_{\text{judge}}$) in Stage 3 results in larger gains by assessing intermediate reasoning steps, particularly enhancing performance on fine-grained localization benchmarks (e.g., +0.8 on OmniBench, +1.3 on MMVU). Finally, incorporating VCRS further stabilizes training by maintaining non-zero advantage baselines via moving averages, allowing the full reward setup to achieve the highest observed performance (e.g., +0.8 on OmniBench, +0.9 on MMVU).
How do training dynamics differ across benchmarks?
The training curves in Figure 6 reveal distinct learning patterns that directly validate our core contributions. Baseline GRPO shows a consistent decline across all benchmarks due to the vanishing advantage problem: when encountering difficult samples where all responses receive similar rewards, the learning signal collapses to zero, stalling further learning. As a result, GRPO quickly plateaus and fails to move beyond solving only the simplest queries.
In contrast, AVATAR shows a dip-and-recovery pattern that reflects the interplay of its key components. The initial drop in performance, especially on challenging benchmarks like OmniBench, Video-Holmes, and MMVU, occurs when the stratified replay buffer transitions from easy to hard samples, exposing the model to its most difficult failures. This performance dip would persist without TAS; however, the U-shaped temporal weighting in TAS provides focused learning signals that extract meaningful gradients from these difficult examples. By emphasizing the planning and synthesis phases that uniform credit assignment would dilute, TAS enables recovery and subsequent improvement. The model’s rebound and higher final performance confirm that our off-policy architecture transforms early failure into learning opportunities through repeated exposure, while TAS ensures each replayed experience yields maximum value. Notably, these gains emerge within just 1000 iterations, underscoring AVATAR’s superior sample efficiency and training effectiveness.
Does AVATAR generalize to smaller models?
To evaluate the generalizability of our approach across different model scales, we conduct experiments with a 3B parameter version of the base Qwen2.5 Omni (Xu et al. 2025) model. Tables 7 and 8 demonstrate that AVATAR consistently outperforms both the baseline and on-policy GRPO across all benchmarks, achieving notable improvements of +3.4 on OmniBench, +2.9 on Video-Holmes, and +2.8 on MMVU. These results validate that AVATAR’s core components, the stratified replay buffer and TAS, effectively enhance multimodal reasoning capabilities regardless of language model scale. The consistent performance patterns across different model sizes further confirm that our off-policy architecture is broadly applicable, with particularly strong benefits observed on complex reasoning tasks where AVATAR’s targeted learning signals prove most impactful for guiding model improvement.
Audio-Visual Benchmarks
Model OmniBench DailyOmni AV-Counting AV-Odyssey WorldSense IntentBench
Baseline 3B Omni 42.4 42.9 19.7 27.2 35.2 59.7
Baseline GRPO 43.1+0.7 43.4+0.5 20.0+0.3 28.5+1.3 35.8+0.6 59.9+0.2
AVATAR 3B 45.8+3.4 44.7+1.8 20.9+1.2 29.8+2.6 37.1+1.9 60.5+0.8
Table 7: Audio-visual benchmarks: 3B model comparison. AVATAR 3B consistently outperforms both the baseline 3B Omni and baseline GRPO across all benchmarks, achieving improvements of +3.4 on OmniBench and +2.6 on AV-Odyssey. The results demonstrate that our off-policy architecture and TAS effectively enhance multimodal reasoning capabilities even with smaller model scales, validating the broad applicability of our approach. Performance improvements over baseline 3B Omni are shown as subscripts.
General Video Understanding Video Reasoning
Model MVBench Video-MME LVBench Video-Holmes MMVU TOMATO
Baseline 3B Omni 59.2 57.6 34.3 39.2 62.0 27.0
Baseline GRPO 59.8+0.6 58.1+0.5 34.7+0.4 40.8+1.6 63.2+1.2 27.3+0.3
AVATAR 3B 61.2+2.0 59.9+2.3 35.8+1.5 42.1+2.9 64.8+2.8 28.4+1.4
Table 8: Video understanding and reasoning benchmarks: 3B model comparison. AVATAR 3B demonstrates superior performance across both general video understanding and complex reasoning tasks, with particularly strong gains on reasoning-heavy benchmarks like Video-Holmes (+2.9) and MMVU (+2.8). The consistent improvements over baseline GRPO highlight how our stratified replay buffer and TAS enhance learning efficiency even with reduced model capacity, confirming that our framework’s benefits scale effectively across different model sizes. Performance improvements over baseline 3B Omni are shown as subscripts.
Hint Generation Prompt for Hard Samples You are an expert AI tutor. Your task is to provide a concise, strategic hint to a student model that is struggling with a difficult audio-visual reasoning problem. The student model is consistently failing this problem and has stopped exploring new reasoning paths. Your hint should guide it out of a local minimum without solving the problem for it. Given the following full problem context: [User Query]: {The original question posed to the student model.} [Audio Caption]: {A transcript of the relevant audio from the video clip.} [Video Context]: {The full video clip associated with the query.} [Ground Truth Answer]: {The correct final answer to the query.} Your Task: Based on the full context and the ground truth, generate a single, short sentence that provides a high-level strategy for how to approach the problem. Constraints: DO NOT reveal or allude to the final answer. DO NOT perform the reasoning steps yourself. Your hint must be a strategic suggestion (e.g., “focus on the audio first,” “compare the objects on the left and right,” “track the object’s movement over time”). Example Hint: first locate the object making the sound, then count Output: Provide ONLY the text of the hint.
Figure 7: The prompt template used to generate guiding hints for the main policy model.
Prompt Templates
To support AVATAR’s training pipeline, we use two prompt templates for different components of our framework. The hint generation prompt (Figure 7) is used by our stratified replay buffer’s hint mechanism when the policy becomes stuck in local optima on challenging samples. This prompt instructs a teacher model to generate strategic hints without revealing the final answer, helping the policy escape stagnation while maintaining the learning challenge. The audio-visual localization judge prompt (Figure 8) powers our stepwise reasoning judge in Stage 3 training by providing granular evaluation criteria across four dimensions: grounding of audio cues, identification of visual objects, spatial localization accuracy, and caption correctness. This offers detailed feedback on cross-modal reasoning quality rather than binary success/failure signals.
Audio-Visual Localization Judge Prompt You are a meticulous and precise Audio-Visual Grounding Evaluator. Your task is to provide a granular, stepwise evaluation of a model’s attempt to localize an object based on audio and visual cues. Given the following inputs, you will score the model’s reasoning and final answer on several criteria. [User Query]: {The user’s original question to the model.} [Audio Caption]: {A transcript of the relevant audio from the video clip.} [Ground Truth Answer]: {The correct description and location of the target object.} [Model’s Reasoning]: {The reasoning process generated by the model inside its <think> tags.} [Model’s Final Answer]: {The final location and description from the model’s <answer> tag.} Your Task: Based on your analysis of all the provided information, you must evaluate the model’s performance on the following four criteria. Provide your judgment as a single JSON object with a score from 0.0 (complete failure) to 1.0 (perfect) for each criterion. 1. Audio Cue Grounding (audio_grounding_score): Did the model’s reasoning correctly identify the key descriptive words in the [Audio Caption] that point to the target? (e.g., did it correctly identify “clapping” as the key sound?). 2. Visual Object Identification (visual_id_score): Based on the audio cue, did the model’s reasoning correctly identify the corresponding visual object in the scene? (e.g., after identifying “clapping,” did it correctly identify the “audience” as the source?). 3. Location Accuracy (location_acc_score): How accurate is the final spatial location (e.g., “Left,” “Center”) provided in the [Model’s Final Answer] when compared to the [Ground Truth Answer]? 4. Caption Correctness (caption_corr_score): How well does the model’s final textual description of the object in its [Model’s Final Answer] match the [Ground Truth Answer]? Output Format: Return ONLY a single JSON object with your scores. Example Output:
{‘‘audio_grounding_score’’: 0.9, ‘‘visual_id_score’’: 1.0, ‘‘location_acc_score’’: 1.0, ‘‘caption_corr_score’’: 0.8}
Figure 8: The prompt template for our VLM-based judge in Stage 3. It guides the judge to produce a granular, stepwise reward by scoring the policy model’s reasoning on audio grounding, visual identification, location accuracy, and caption correctness.
Qualitative Analysis
Figure 9 illustrates the qualitative differences between baseline GRPO and AVATAR on audio-visual reasoning tasks. AVATAR demonstrates superior cross-modal integration by linking visual cues (“tense expression, his eyes darting around”) with audio analysis (“hurried and tense tone when he speaks”), while baseline GRPO makes disconnected observations, such as “He looks worried which indicates stress” without linking modalities. The responses highlight the effectiveness of TAS, evident in the model’s ability to plan by establishing comprehensive scene context, focus on relevant cues while avoiding redundant pattern matching (e.g., “Black clothing often symbolizes negativity”), and synthesize complex emotional dynamics rather than relying on surface-level assumptions (“something bad might be happening”). AVATAR also shows improved temporal reasoning by tracking emotional progression (“tone shifting from calm to anxious”) and enhanced contextual understanding through precise dialogue interpretation (“Sorry, I have a train to catch” indicating abrupt departure). In contrast, baseline GRPO tends to fall back on generic genre classifications (“suggests a thriller or drama genre”). These qualitative differences validate our framework’s ability to enhance multimodal reasoning quality through targeted credit assignment and structured learning signals.
Figure 9: Qualitative comparison between Baseline GRPO and AVATAR. AVATAR demonstrates better cross-modal integration and structured reasoning, while Baseline GRPO shows surface-level observations and disconnected audio-visual analysis. Green highlighting shows AVATAR’s effective synthesis of multimodal cues, red highlighting reveals GRPO’s uniform credit assignment limitations.
1def compute_avatar_loss(policy_logprobs: torch.Tensor,
2 old_logprobs: torch.Tensor,
3 advantages: torch.Tensor,
4 completion_mask: torch.Tensor,
5 lambda_tas: float,
6 epsilon: float) -> torch.Tensor:
7
8 # 1. Standard GRPO Clipped Surrogate Objective
9 # ---------------------------------------------
10 # Calculate the importance sampling ratio per token
11 ratio = torch.exp(policy_logprobs - old_logprobs)
12
13 # Clip the ratio to prevent excessively large policy updates
14 clipped_ratio = torch.clamp(ratio, 1 - epsilon, 1 + epsilon)
15
16 # The advantage is broadcasted to match the shape of the per-token ratios
17 advantages_per_token = advantages.unsqueeze(1)
18
19 # Calculate the unclipped and clipped surrogate losses per token
20 loss_unclipped = ratio * advantages_per_token
21 loss_clipped = clipped_ratio * advantages_per_token
22
23 # The standard per-token loss is the minimum of the two, negated
24 per_token_loss = -torch.min(loss_unclipped, loss_clipped)
25
26 # 2. Temporal Advantage Shaping (TAS)
27 # ------------------------------------
28 # Create a tensor to hold the temporal weights, initialized to a baseline of 1.0
29 temporal_weights = torch.ones_like(per_token_loss)
30
31 # Calculate U-shaped weights for each sequence in the batch
32 for i in range(per_token_loss.size(0)):
33 seq_len = completion_mask[i].sum().item()
34 if seq_len > 1:
35 # Normalize token positions to a range of [0, 1]
36 positions = torch.arange(seq_len, device=per_token_loss.device, dtype=torch.float32)
37 normalized_positions = positions / (seq_len - 1)
38
39 # U-Shaped Curve: y = 1.0 + lambda * (2x - 1)^2
40 parabolic_curve = (2 * normalized_positions - 1)**2
41 u_curve_weights = 1.0 + (lambda_tas * parabolic_curve)
42
43 # Apply these weights to the relevant part of the sequence
44 temporal_weights[i, :seq_len] = u_curve_weights
45
46 # 3. Final Loss Calculation
47 # --------------------------
48 # Apply the temporal weights to shape the credit assignment
49 shaped_per_token_loss = per_token_loss * temporal_weights
50
51 # Apply the completion mask to only consider the reasoning tokens
52 masked_loss = shaped_per_token_loss * completion_mask
53
54 # Aggregate the loss for the entire batch
55 final_loss = masked_loss.sum() / completion_mask.sum().clamp(min=1.0)
56
57 return final_loss
Listing 1: Pseudocode for TAS.